OpenStudy (anonymous):

x2 + 2x what value completes the square for each expression?

OpenStudy (anonymous):

x^2 +2x+(2/2)^2 = 0 +(2/2)^2 x^2 +2x +1^2 = 0+1^2 x^2 +2x +1 = 0+1 (x+1)^2 = 1 (x+1)^2 - 1 = 0